# Stochastic Path Planning in Correlated Obstacle Fields

### v1

@@ 5.1 Identifying the Decision Set with Search Space Reduction
5.1.1 Macro-Path and Decision Set Construction
Each decision $d_{t}\in\mathcal{D}_{t}(S_{t})$ moves the agent from $v_{t}$ to a location $v_{t+1}$ via a macro-path segment. Following the sensing and disambiguation (if needed), the state transitions from $S_{t}$ to $S_{t+1}$. The key question is: which ambiguous obstacles should be prioritized for disambiguation and sensing, and at which adjacent vertices (referred to as the disambiguation vertex) should this be implemented? We use an optimistic greedy approach that iteratively identifies candidates to be included in $\mathcal{D}_{t}(S_{t})$.
@@ 13    $\mathcal{D}_{t}\leftarrow\mathcal{D}_{t}\cup\{v(x_{i})\}$;
While including more candidates provides more information to enhance the decision-making process, it may lead to over-exploration and waste of computational resources (MacDonald and Smith,, 2020). The following proposition (Proposition 5.1) ensures that our approach (Algorithm 1) effectively avoids over-exploration and stops when no further improvement can be achieved over a pure exploitation decision (i.e., the best path avoiding all ambiguous and true obstacles based on current knowledge). Define a candidate in the decision set $\mathcal{D}_{t}$ to be a valid candidate if it can potentially achieve improvement over pure exploitation decision, we show that:
Proposition 5.1.
The decision set $\mathcal{D}_{t}$ determined by the iterative optimistic greedy approach (Algorithm 1) contains only valid candidates, either disambiguation vertices or the goal $g$.
Proof.
At each iteration, we find the minimum cost path under optimism assumption and identify the first ambiguous obstacle $x_{i}$ along this path. The cost of this path can be decomposed as:
$\ell_{free}(v_{t},x_{i})+c(x_{i})+\mathcal{C}_{optimism}(x_{i},g),$
where $\ell_{free}(v_{t},x_{i})$ represents the length of obstacle-free path from current vertex $v_{t}$ to the obstacle $x_{i}$, $c(x_{i})$ is obstacle’s disambiguation cost and $\mathcal{C}_{optimism}(x_{i},g)$ is the minimum total traversal cost from obstacle $x$ to $g$ by assuming all obstacles of uncertain status are false (i.e., traversable). Since the algorithm stops when $g$ is added to the set, it is equivalent to adding disambiguation vertices of all obstacles satisfying the following condition,
$\ell_{free}(v_{t},g)\geq\ell_{free}(v_{t},x)+c(x)+\mathcal{C}_{optimism}(x,g).$
This indicates that disambiguating $x$ could potentially yield a cheaper path than the best obstacle-free path (i.e., the pure exploitation path). Otherwise, $x$ will not be added to $\mathcal{D}_{t}$ if its best possible performance cannot beat the best obstacle-free path. ∎
@@ 5.1.2 Search Space Reduction
To further improve efficiency, we reduce the search space by discarding obstacles and their associated paths whose best performance cannot exceeding the current best solution. Combining the decision set construction with search space reduction strategy, we minimize computation cost while preserving only the most promising traversal options. This enables the optimal policy to scale effectively to larger and more complex environments.
@@ 6.1 Optimistic Policy Iteration with Information Integration
Policy iteration (PI) is a foundamental framework in RL (Sutton and Barto,, 2018). Its core idea is the alternating implementation between two key steps: policy evaluation and policy improvement. The policy evaluation step involves generating interaction experience under a fixed policy. The resulting observations are used to update the state value functions. In the policy improvement step, the updated value function estimates are used to define a better policy, supporting further refinement. This alternating process drives the estimated value functions $V(s)$ towards their optimal values $V^{*}(s)$.
@@ 6.2 Model-Based and Model-Free Reinforcement Learning Structure
Within this information-guided OPI framework, we integrate model-based and model-free components. During policy evaluation, the model-based component uses posterior sampling to generate probabilistic environment models that guide exploration. The model-free component complements this by directly learning value functions from simulated experiences, refining estimates while adapting to sequentially updated posterior distributions. The interplay of these two components creates a robust learning environment, effectively addressing uncertainty in obstacle status (i.e., blockage probabilities).
@@
Our approach aligns with Dyna-based methods in hybrid-RL literature (Sutton,, 1991; Silver et al.,, 2008), where the model-based component provides dynamic information to enhance model-free learning and improve sampling efficiency. By incorporating posterior sampling through Bayesian updating, we emphasize a systematic approach for handling uncertainty and optimizing exploration-exploitation trade-off, resulting in more robust base policy learning. The steps are detailed in Algorithm 4.
@@ 6.3.2 Distributional RL Approach
Otherwise, we directly insert $\mathcal{C}^{\scriptscriptstyle IG}_{i}(s)$ into the support set and assign $\alpha_{j+1}=1$, since it is the first observation of that value. If $\mathcal{C}^{\scriptscriptstyle IG}_{i}(s)$ is already included in the support, we update using the Bayesian update rule. This refinement guarantees that the support of $V(s)$ aligns with the actual values which can be observed, enhancing the accuracy of distribution representation.
@@ 6.4 Offline Learning Framework Properties
In this section, we establish theoretical results to justify our offline learning strategy. First, we demonstrate the benefit of incorporating information gain into value function updating process, then show that the information gain should maintain diminishing impact on the decision making due to submodularity property. We also prove the convergence of OPI under posterior sampling.
Theorem 6.1.
Given that the state space and the decision set are finite in our problem setting, assume every policy $\pi$ is proper (i.e., reaches the goal with probability 1 in finitely many steps), and the immediate cost $\mathcal{C}(s,d)$ is bounded, $\sup_{s,d}\mathcal{C}(s,d)<\infty$. Define two Bellman operators
$(\mathbf{T}V)(s)=\min_{d\in\mathcal{D}(s)}\bigl\{\mathcal{C}(s,d)+\mathbf{E}[V(s^{\prime})\mid s,d]-G(d)\bigr\},\;(\mathbf{T}_{\mathrm{b}}V)(s)=\min_{d\in\mathcal{D}(s)}\bigl\{\mathcal{C}(s,d)+\mathbf{E}[V(s^{\prime})\mid s,d]\bigr\}.\$
Let $V^{\!*}$ and $V^{b}$ denote their fixed points. Then:
@@
    Let $\pi^{b}$ be the optimal policy obtained from $V^{b}$. Then for any start state $s_{0}$,
    $V^{b}(s_{0})-V^{*}(s_{0})\geq\mathbf{E}_{\pi^{b}}\left[\sum_{t=0}^{T-1}G(d_{t})\right],$
    where the expectation is taken over trajectory distribution induced by $\pi^{b}$ and transition probabilities.
@@
Because $G(d)\geq 0$, we have point-wise operator ordering $(\mathbf{T}V)(s)\leq(\mathbf{T}_{\mathrm{b}}V)(s)$ for every bounded $V$ and state $s$. Bellman operators are monotone contractions, hence iteration from any common initial value preserves this order, leading to $V^{\!*}\leq V^{b}$.
Given a state $s$, denote
$d^{b}(s)=\pi^{b}(s),\quad\Delta(s)=V^{b}(s)-V^{\!*}(s)\;\;\geq 0.$
Because $V^{\!*}$ is the minimum in $\mathbf{T}$, it satisfies, for any decisionădd,
$V^{\!*}(s)\leq\mathcal{C}(s,d)+\mathbf{E}[V^{\!*}(s^{\prime})\mid s,d]-G(d),$
where $s^{\prime}$ is the next random state given current state-decision pair. Let $d=d^{b}$ and subtracting the equation for $V^{b}$ results in the inequality:
$\Delta(s)\;\geq\;G(d^{b})\;+\;\mathbf{E}\bigl[\Delta(s^{\prime})\mid s,d^{b}\bigr].$
Cumulatively, it leads to
$\Delta(s_{0})\;\geq\;\sum_{t=0}^{T-1}G(d^{b}_{t}),$
taking expectation with respect to the trajectory distributions induced by $\pi^{b}$ and transition probabilities yields the desired result. ∎
Proposition 6.2 (Submodularity of information gain).
Consider the log-odds vector $Y\sim\mathcal{N}(0,K)$ in our SCOS model. Let $\tilde{Y}_{X^{\prime}}$ be the random vector corresponding to potential log-odds observations of obstacle set $X^{\prime}\subseteq X$. In the context of observations with Gaussian noise, define the information gain as the mutual information
$\mathbf{I}(\tilde{Y}_{X^{\prime}})=\text{MI}(Y_{X^{\prime}};\tilde{Y}_{X^{\prime}})=\tfrac{1}{2}\log\det\!\big(I+\Sigma_{X^{\prime}}^{-1}K_{X^{\prime}X^{\prime}}\big),$
where $K_{X^{\prime}X^{\prime}}$ is the covariance matrix corresponding to the obstacle set $X^{\prime}$, and $\Sigma_{X^{\prime}}$ is the diagonal matrix containing observation noise variance. Then:

### v2

@@ 5.1 Identifying the Decision Set with Search Space Reduction
5.1.1 Decision Structure
At decision step $t$, the decision set including all possible decisions (i.e., stopping vertices can be reached via obstacle-free path from $v_{t}$) is
[…]
Implementing $d_{t}\in\mathcal{D}_{t}(S_{t})$ moves the agent from $v_{t}$ to the chosen stopping vertex, incurring the segment length. If the stopping vertex is adjacent to ambiguous obstacles, the agent may disambiguate at additional costs, and collect readings of obstacles within the sensing range.
5.1.2 Optimistic Candidate Generation and Soundness
To construct $\mathcal{D}_{t}$, two key questions should be addressed: which ambiguous obstacles should be prioritized for disambiguation and sensing, and at which adjacent vertices (referred to as the disambiguation vertex) should this be implemented? We use an optimistic greedy approach that iteratively identifies candidates.
@@ 13    $\mathcal{D}_{t}\leftarrow\mathcal{D}_{t}\cup\{v(x_{i})\}$;
This iterative optimistic-greedy procedure adds only candidates with potential improvement over a pure exploitation decision, therefore avoiding over-exploration and wasted computation (MacDonald and Smith,, 2020). The formal soundness guarantee is established in the Proposition 5.6 (Section 5.1.4).
@@ 5.1.3 Search Space Reduction
To further improve efficiency, we reduce the search space by discarding obstacles and their associated paths whose best performance cannot exceed the current best solution. Combining the decision set construction with search space reduction strategy, we minimize computation cost while preserving only the most promising traversal options. This enables the optimal policy to scale effectively to larger and more complex environments.
@@ 11    $X_{\text{discard}}\leftarrow X_{\text{discard}}\cup\{x\}$;
5.1.4 Soundness of the Decision Set
We work on a finite, undirected graph $G=(V,E)$ with strictly positive edge lengths. Let $v_{t}\in V$ be the current vertex and $g\in V$ the goal. Let $X^{O}$ denote the set of (known) true obstacles and $X^{U}$ the set of ambiguous obstacles. For each ambiguous obstacle $x\in X^{U}$, fix a deterministic tie-breaking rule that selects a disambiguation vertex $v(x)$ whenever multiple choices are possible (e.g., the vertex encountered along the chosen optimistic shortest path).
Definition 5.1 (Pure exploitation baseline).
Let $\mathcal{C}_{\mathrm{exploit}}(v_{t})$ be the length of a shortest path from $v_{t}$ to $g$ that avoids all true and ambiguous obstacles (i.e., a path feasible with respect to $X^{O}\cup X^{U}$); set $\mathcal{C}_{\mathrm{exploit}}(v_{t})=+\infty$ if no such path exists.
Definition 5.2 (Per-obstacle optimistic lower bound).
For $x\in X^{U}$, define
$\underline{\mathcal{C}}(x)\ :=\ \ell_{\mathrm{free}}(v_{t},x)\ +\ c(x)\ +\ \mathcal{C}_{\mathrm{optimism}}(x,g),$
where $\ell_{\mathrm{free}}(v_{t},x)$ is the length of a shortest obstacle-free path segment from $v_{t}$ to the disambiguation vertex $v(x)$ (i.e., a segment that intersects neither true nor ambiguous obstacles), $c(x)$ is the disambiguation cost at $x$, and $\mathcal{C}_{\mathrm{optimism}}(x,g)$ is the shortest-path cost from $v(x)$ to $g$ under the optimism assumption that all ambiguous obstacles are traversable (true obstacles remain forbidden).
Definition 5.3 (Optimistic evaluation and minimizer).
For any $v\in V$, define the optimistic evaluation of starting at $v$ by
$\mathrm{OptEval}(v)\ :=\ \min\!\Big\{\,\mathcal{C}_{\mathrm{optimism}}(v,g)\,,\ \min_{x\in X^{U}}\ \underline{\mathcal{C}}(x)\,\Big\}.$
Equivalently: either reach $g$ without touching any ambiguous obstacle (first term), or first disambiguate some $x$ and then proceed optimistically (second term). An optimistic-evaluated minimizer is any path that attains $\mathrm{OptEval}(v_{t})$ and, if it touches an ambiguous obstacle, let $x$ denote the first such obstacle on that path.
Lemma 5.4 (Optimistic evaluation decomposition).
Let $P^{\star}$ be an optimistic-evaluated minimizer from $v_{t}$. If $P^{\star}$ reaches $g$ without meeting any ambiguous obstacle, then $\mathrm{OptEval}(v_{t})=\mathcal{C}_{\mathrm{optimism}}(v_{t},g)$. Otherwise, if $x$ is the first ambiguous obstacle met along $P^{\star}$, then
$\mathrm{OptEval}(v_{t})\ =\ \ell_{\mathrm{free}}(v_{t},x)\ +\ c(x)\ +\ \mathcal{C}_{\mathrm{optimism}}(x,g)\ =\ \underline{\mathcal{C}}(x).$
Proof.
By definition of $\mathrm{OptEval}$, either the no-ambiguity option is optimal, or some $x$ is optimal to disambiguate first. In the latter case, the prefix to $v(x)$ is obstacle-free by definition of “first ambiguous,” so its length is $\ell_{\mathrm{free}}(v_{t},x)$; at $v(x)$ we incur $c(x)$; and the suffix from $v(x)$ to $g$ is a shortest path under optimism (optimal substructure of shortest paths with positive edge lengths). Concatenation yields $\underline{\mathcal{C}}(x)$. ∎
Lemma 5.5 (Optimistic evaluation vs. exploitation).
We have
$\mathrm{OptEval}(v_{t})\ \leq\ \mathcal{C}_{\mathrm{exploit}}(v_{t}).$
Proof.
Consider an exploitation-shortest path (Definition 5.1): it avoids all ambiguous obstacles, hence it is also feasible under optimism, and its optimistic evaluation equals its length $\mathcal{C}_{\mathrm{exploit}}(v_{t})$ (no $c(x)$ is incurred). Taking the minimum over all options in $\mathrm{OptEval}(v_{t})$ cannot exceed this value. ∎
Proposition 5.6 (Soundness of the optimistic candidate set).
Let $\mathcal{D}_{t}$ be the decision set built by Algorithm 1. Then every obstacle $x$ whose disambiguation vertex is added to $\mathcal{D}_{t}$ satisfies
$\underline{\mathcal{C}}(x)\ \leq\ \mathcal{C}_{\mathrm{exploit}}(v_{t}),$
and, if $g$ is added, it is trivially valid. In particular,
$\mathcal{D}_{t}\ \subseteq\ \{\,g\,\}\ \cup\ \bigl\{\,v(x)\ :\ x\in X^{U},\ \underline{\mathcal{C}}(x)\leq\mathcal{C}_{\mathrm{exploit}}(v_{t})\,\bigr\}.$
Proof.
By Lemma 5.4, in any iteration that adds a disambiguation vertex for obstacle $x$, the evaluation value equals $\underline{\mathcal{C}}(x)$. By Lemma 5.5, $\mathrm{OptEval}(v_{t})\leq\mathcal{C}_{\mathrm{exploit}}(v_{t})$, hence $\underline{\mathcal{C}}(x)\leq C_{\mathrm{exploit}}(v_{t})$ for that $x$. If the minimizer reaches $g$ without touching an ambiguous obstacle, we add $g$ and terminate. Repeating the argument per iteration yields the stated inclusion. ∎
The test $\underline{\mathcal{C}}(x)\leq\mathcal{C}_{\mathrm{exploit}}(v_{t})$ formalizes “potential improvement over pure exploitation.” Proposition 5.6 shows soundness (no invalid candidates are added). The procedure is generally not complete: there may exist obstacles with $\underline{\mathcal{C}}(x)\leq\mathcal{C}_{\mathrm{exploit}}(v_{t})$ that are not the first ambiguous obstacle on any optimistic-evaluated minimizer and hence are not added.
Strictly positive edge lengths preclude zero-cost cycles and ensure shortest-path problems are well-defined (ties broken deterministically). If $\mathcal{C}_{\mathrm{exploit}}(v_{t})=+\infty$ (no exploitation path exists), the inequality $\underline{\mathcal{C}}(x)\leq\mathcal{C}_{\mathrm{exploit}}(v_{t})$ holds vacuously for any $x$ with finite $\underline{\mathcal{C}}(x)$, and the proof remains valid.
@@ 6.1 Optimistic Policy Iteration with Information Integration
Policy iteration (PI) is a fundamental framework in RL (Sutton and Barto,, 2018). Its core idea is the alternating implementation between two key steps: policy evaluation and policy improvement. The policy evaluation step involves generating interaction experience under a fixed policy. The resulting observations are used to update the state value functions. In the policy improvement step, the updated value function estimates are used to define a better policy, supporting further refinement. This alternating process drives the estimated value functions $V(s)$ towards their optimal values $V^{*}(s)$.
@@ 6.2 Model-Based and Model-Free Reinforcement Learning Structure
Within this information-guided OPI framework, we adopt a hybrid structure. During policy evaluation, the model-based component uses posterior sampling to generate probabilistic environment models that guide exploration. The model-free component complements this by directly learning value functions from simulated experiences, refining estimates while adapting to sequentially updated posterior distributions. The interplay of these two components creates a robust learning environment, effectively addressing uncertainty in obstacle status (i.e., blockage probabilities).
@@
Our approach aligns with Dyna-based hybrid RL (Sutton,, 1991; Silver et al.,, 2008), where the model-based component provides dynamic information to enhance model-free learning and improve sampling efficiency. By incorporating posterior sampling through Bayesian updating, we emphasize a systematic approach for handling uncertainty and optimizing exploration-exploitation trade-off, resulting in more robust base policy learning. The steps are detailed in Algorithm 4.
@@ 6.3.2 Distributional RL Approach
Otherwise, we directly insert $\mathcal{C}^{\scriptscriptstyle IG}_{i}(s)$ into the support set and assign $\alpha_{j+1}=1$, since it is the first observation of that value. If $\mathcal{C}^{\scriptscriptstyle IG}_{i}(s)$ is already included in the support, we update using the Bayesian update rule. This refinement guarantees that the support of $V(s)$ aligns with the actual values which can be observed, enhancing the accuracy of distribution representation. Additionally, this support refinement procedure maintains stability in the mean estimate, with bounded changes that shrink as more observations accumulate, which is formally established in Lemma 6.6 (Section 6.4).
The theoretical analysis of distributional RL to our approximate Bayesian setting, and evaluation of distributional calibration (e.g., prediction interval coverage, CRPS) remains an important direction for future investigation.
@@ 6.4 Offline Learning Framework Properties
In this section, we establish theoretical results to justify our offline learning strategy. First, we demonstrate the benefit of incorporating information gain into value function updating process, then show that the information gain should maintain diminishing impact on the decision making due to submodularity property. We also prove the convergence of OPI under posterior sampling, and show that our support refinement procedure for distributional learning maintains stability with bounded, diminishing changes in mean estimates.
6.4.1 Benefit of Nonnegative Bonus Shaping
We consider an undiscounted, episodic stochastic shortest path (SSP) MDP with finite state set $\mathcal{S}$ and finite feasible decision sets $\mathcal{D}(s)$. There is an absorbing goal $g\in\mathcal{S}$ with $\mathcal{C}(g,d)=0$ and $P(g\mid g,d)=1$ for all $d\in\mathcal{D}(g)$. A policy $\pi$ is proper if the hitting time $\tau:=\inf\{t\geq 0:S_{t}=g\}$ is finite a.s. from any start state.
Assume:
  1. (A1)
    All admissible policies are proper, and base one–stage costs are nonnegative and bounded: $0\leq\mathcal{C}(s,d)\leq\overline{\mathcal{C}}<\infty$.
  2. (A2)
    The (exploration) bonus is bounded and nonnegative: $G:\{(s,d):d\in\mathcal{D}(s)\}\to[0,\infty)$ with $\sup_{s,d}G(s,d)<\infty$.
For a (history-dependent) policy $\pi$, define the base and bonus-shaped performances from $s$:
$\small J^{\pi}_{\mathrm{b}}(s):=\mathbb{E}_{\pi}\!\Big[\textstyle\sum_{t=0}^{\tau-1}\mathcal{C}(S_{t},D_{t})\ \Big|\ S_{0}=s\Big],\quad J^{\pi}_{*}(s):=\mathbb{E}_{\pi}\!\Big[\textstyle\sum_{t=0}^{\tau-1}\big(\mathcal{C}(S_{t},D_{t})-G(S_{t},D_{t})\big)\ \Big|\ S_{0}=s\Big].$}
Let $J^{\mathrm{b}}(s):=\inf_{\pi}J^{\pi}_{\mathrm{b}}(s)$ and $J^{*}(s):=\inf_{\pi}J^{\pi}_{*}(s)$. (Under (A1)(A2), $J^{\mathrm{b}}$ is finite and equals the minimal nonnegative solution of the SSP optimality equations. The shaped value $J^{*}$ is well defined as the optimal value of the shaped problem; the ordering $J^{*}\leq J^{\mathrm{b}}$ in Theorem 6.1 does not require the shaped costs to be nonnegative.)
Theorem 6.1 (Benefit of nonnegative bonus).
Under (A1)(A2):
@@ 6.4.1 Benefit of Nonnegative Bonus Shaping
    Let $\pi^{\mathrm{b}}$ be an optimal policy for the base problem, i.e., $J^{\mathrm{b}}(s)=J^{\pi^{\mathrm{b}}}_{\mathrm{b}}(s)$. Then for any start state $s_{0}$,
    $J^{\mathrm{b}}(s_{0})-J^{*}(s_{0})\;\;\geq\;\;\mathbb{E}_{\pi^{\mathrm{b}}}\!\Big[\textstyle\sum_{t=0}^{\tau-1}G(S_{t},D_{t})\ \Big|\ S_{0}=s_{0}\Big].$
@@
Step 1 (policywise identity). For any fixed policy $\pi$ and any start state $s$,
$J^{\pi}_{\mathrm{b}}(s)-J^{\pi}_{*}(s)=\mathbb{E}_{\pi}\!\Big[\textstyle\sum_{t=0}^{\tau-1}G(S_{t},D_{t})\ \Big|\ S_{0}=s\Big],$
which follows directly from the definitions by linearity of expectation.
Step 2 (ordering of optimal values). By Step 1, since $G\geq 0$, $J^{\pi}_{*}(s)\leq J^{\pi}_{\mathrm{b}}(s)$ for all $\pi$ and $s$. Taking infimum over $\pi$ on both sides yields $J^{*}(s)\leq J^{\mathrm{b}}(s)$, proving (i).
Step 3 (gap bound at a base-optimal policy). For the base-optimal policy $\pi^{\mathrm{b}}$,
$J^{\mathrm{b}}(s_{0})-J^{*}(s_{0})\;\;\geq\;\;J^{\pi^{\mathrm{b}}}_{\mathrm{b}}(s_{0})-J^{\pi^{\mathrm{b}}}_{*}(s_{0})\;=\;\mathbb{E}_{\pi^{\mathrm{b}}}\!\Big[\textstyle\sum_{t=0}^{\tau-1}G(S_{t},D_{t})\ \Big|\ S_{0}=s_{0}\Big],$
where the inequality uses $J^{*}(s_{0})\leq J^{\pi^{\mathrm{b}}}_{*}(s_{0})$ (optimality of $J^{*}$), and the equality is Step 1. This proves (ii). ∎
Remark 6.2 (Bellman-operator view).
Define the dynamic-programming operators on bounded $V:\mathcal{S}\to\mathbb{R}$ by
$(\mathbf{T}_{*}V)(s)$ $:=\min_{d\in\mathcal{D}(s)}\Big\{\mathcal{C}(s,d)-G(s,d)+\sum_{s^{\prime}}P(s^{\prime}|s,d)V(s^{\prime})\Big\},$
$(\mathbf{T}_{\mathrm{b}}V)(s)$ $:=\min_{d\in\mathcal{D}(s)}\Big\{\mathcal{C}(s,d)+\sum_{s^{\prime}}P(s^{\prime}|s,d)V(s^{\prime})\Big\}.$
Then $(\mathbf{T}_{*}V)(s)\leq(\mathbf{T}_{\mathrm{b}}V)(s)$ for all $V,s$ because $G\geq 0$. Under SSP with nonnegative base costs, the minimal nonnegative fixed point of $\mathbf{T}_{\mathrm{b}}$ is $J^{\mathrm{b}}$. If, in addition, $G(s,d)\leq\mathcal{C}(s,d)$ for all $(s,d)$, then $\mathbf{T}_{*}$ also has a minimal nonnegative fixed point equal to $J^{*}$; otherwise $J^{*}$ should be interpreted as the optimal value solving the shaped problem without the nonnegativity qualifier. The ordering $J^{*}\leq J^{\mathrm{b}}$ follows from the pointwise operator inequality and a standard telescoping argument along trajectories until absorption.
For a fixed finite horizon $T<\infty$, replace $\tau$ by $T$ throughout; all conclusions remain valid without additional assumptions.
6.4.2 Submodularity of Information Gain
Linear–Gaussian Sensing Model.
Let $Y=(Y_{x})_{x\in X}\sim\mathcal{N}(0,K)$ with $K\succ 0$. For any $A\subseteq X$, define noisy observations
$\widetilde{Y}_{A}\ :=\ Y_{A}+\varepsilon_{A},\qquad\varepsilon_{A}\sim\mathcal{N}(0,\Sigma_{A}),$
with $\varepsilon_{A}$ independent of $Y$ and of $\varepsilon_{B}$ for disjoint $A,B$. Assume throughout the independent-noise case $\Sigma_{A}=\mathrm{diag}\big((\sigma_{x}^{2})_{x\in A}\big)\succ 0$. Define the set function
$\mathbf{I}(A)\ :=\ \mathrm{MI}\big(Y_{A};\ \widetilde{Y}_{A}\big)\ =\ \tfrac{1}{2}\log\det\!\big(I+\Sigma_{A}^{-1/2}K_{AA}\Sigma_{A}^{-1/2}\big)\ =\ \tfrac{1}{2}\bigl[\log\det(K_{AA}+\Sigma_{A})-\log\det(\Sigma_{A})\bigr].$
Proposition 6.3.
Under the model above, $\mathbf{I}:2^{X}\to\mathbb{R}_{\geq 0}$ satisfies: